OpenStudy (anonymous):

I need help finishing a problem I am almost done with it but I got stuck!!! Here is the question. Using the Law of Cosine, in triangle RST, if r=14, s=9, t= 9 find the angle of S. I am down to -151/-168=cos S but when I divide that it doesn't work out!!!

OpenStudy (anonymous):

@hartnn can you help me?

hartnn (hartnn):

how did u get -151?

hartnn (hartnn):

and 168 too :P which formula did you use? :O

OpenStudy (anonymous):

This is what I did can you tell me if my answer is right?

hartnn (hartnn):

ohhhhhh, your t= 6 ??? you put t=9 in your question

OpenStudy (anonymous):

Ohhh opps sorry, I wasn't looking at the question when I typed it.

hartnn (hartnn):

yes, your work is good angle S is indeed 26 degrees :)

OpenStudy (anonymous):

Thank you. I have more so when I finish them can I post them to see if they are right?

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

@hartnn Is this right?

hartnn (hartnn):

that is also correct :) triangle sides are 15,28,17, right ? then yes, angle is 27 :)

OpenStudy (anonymous):

Okay thank you, haha one more. :)

hartnn (hartnn):

bring it on!

OpenStudy (anonymous):

I need help on this one the question is Using the Law of Sine, in triangle ABC, if angle c= 135, angle b=10, side c= 45, what is the measure of a?

hartnn (hartnn):

find measure of angle a, first! know how to ?

hartnn (hartnn):

C=135, B=10, A= find out c= 45, a= required.

hartnn (hartnn):

A+B+C= 180

OpenStudy (anonymous):

dont you just add up the angles you know then subtract them from 180

OpenStudy (anonymous):

a=25

OpenStudy (anonymous):

i mean 35 sorry

hartnn (hartnn):

yes, A =35 (sorry i lost my internet connection) now use the sine law ! \(\huge \dfrac{a}{\sin A }=\dfrac{c}{\sin C}\) only 'a' is unknown! :)

OpenStudy (anonymous):

this is what i did

hartnn (hartnn):

but you already know angle A .-. you need side =a so plug in A =35, C =135, c =45 in the sine law formula and find 'a' :)

OpenStudy (anonymous):

I'm confused on how to do that

hartnn (hartnn):

\(\huge \dfrac{a}{\sin 35}=\dfrac{45}{\sin 135}\)

hartnn (hartnn):

multiply sin 35 on both sides

OpenStudy (anonymous):

I'm a little confused

hartnn (hartnn):

\(\huge \dfrac{a}{\sin 35}\times \sin 35=\dfrac{45}{\sin 135}\times \sin 35 \\ ~\\ \huge a=\dfrac{45}{\sin 135}\times \sin 35 =... \)

hartnn (hartnn):

if you're still confused, use a calculator :)

OpenStudy (anonymous):

I can't see all of the equation you posted

hartnn (hartnn):

:O you can't ? ... ok, a/ sin 35 = 45 /sin 135 a = 45 sin 35 /sin 135 thats what i wrote!

OpenStudy (anonymous):

ohh okay thank you

hartnn (hartnn):

^_^

OpenStudy (anonymous):

I got 34.2

hartnn (hartnn):

my calculator gave me 36.5

OpenStudy (anonymous):

i did it wrong the first time the second time i got 36.5

hartnn (hartnn):

cool! :)

OpenStudy (anonymous):

Thank you so much!!! :)

hartnn (hartnn):

welcome ^_^